OpenStudy (anonymous):

What is the quotient: (3x^3 + 14x^2 + 13x - 6) / (3x - 1) ?

OpenStudy (abdullahm):

Do you know group factoring?

OpenStudy (anonymous):

No

OpenStudy (abdullahm):

Group factoring is like: \[x^3 + x^2 + x + 1 = x^2(x+1) + 1(x+1) = (x+1)(x^2+1)\]

OpenStudy (anonymous):

I don't really remember learning about that, but I think I understand a little bit

OpenStudy (abdullahm):

so you need to group factor the numerator

OpenStudy (abdullahm):

oh, wait. It can't be group factored >.<

OpenStudy (abdullahm):

The Rational Root Theorem tells you that if the polynomial has a rational zero then it must be a fraction p/q, where p is a factor of the trailing constant and q is a factor of the leading coefficient. and so using that, these could be possible roots \[\pm \frac{ 1 }{ 1 } , ~ \pm \frac{ 1 }{ 3 } , ~ \pm \frac{ 2 }{ 1 } , ~ \pm \frac{ 2 }{ 3 } , ~ \pm \frac{ 3 }{ 1 } , ~ \pm \frac{ 3 }{ 3 } , ~ \pm \frac{ 6 }{ 1 } , ~ \pm \frac{ 6 }{ 3 } ~\]

OpenStudy (abdullahm):

solve 3x-1 = 0 x = 1/3 and since that is one of the possible roots, we can use that since it might be divisible by it.

OpenStudy (abdullahm):

can someone explain? ;-; this is hard to show online compared to in-person with a piece of paper ;-;

OpenStudy (abdullahm):

someone else*

Nnesha (nnesha):

long division

OpenStudy (abdullahm):

ik, but latex isn't that easy to use to show that ;-; and the draw tool isn't that easy either ;-;

Nnesha (nnesha):

for long division 5 steps) \(\huge\color{green}{{1:}}\) Divide the first terms \(\huge\color{green}{{2:}}\) multiply(distribute) \(\huge\color{green}{{3:}}\) subtract all terms \(\huge\color{green}{{4:}}\) carry down \(\huge\color{green}{{5:}}\) repeat

Nnesha (nnesha):

haha true that !

Nnesha (nnesha):

 divide first terms \[\huge\rm \frac{ 3x^3 }{ 3x } = ???\] then multiply the divisor with that